# The Cohen–Lenstra moments over function fields via the stable homology of non-splitting Hurwitz spaces

### v2

@@ A.4. A simple topological model
We finish the proof by showing that show that $\operatorname{colim}_{\alpha\in P}\pi_{0}\Omega X^{\prime}_{\alpha}\to\operatorname{colim}_{\alpha\in P}\pi_{0}\Omega X_{\alpha}$ is a bijection. We first check it is surjective. Given a point $*\in\Omega X_{\alpha}$, taking $K$ to be the compact space $*_{+}$ and $K^{\prime}\subset K$ the base point, the hypotheses applied to the pointed map $*_{+}\to\Omega X_{\alpha}$ shows surjectivity. For injectivity, if $[0,1]_{+}\to\Omega X_{\alpha}$ is a path between two points with the end points factoring through $\Omega X^{\prime}_{\alpha}$, then applying the hypothesis to this map, we get a map $([0,1]\times[0,1])_{+}\to\Omega X_{\beta}$, and the part of the boundary $\{0,1\}\times[0,1]\cup[0,1]\times\{1\}$ yields a path in $\Omega X^{\prime}_{\beta}$ between the two points. ∎
Construction A.4.0.
Choose a smooth function $f:[0,1]\to\mathbb{R}$ such that $f([0,\frac{1}{2}))<0$ and $f(1-x)=-f(x)$. In particular, $f(1/2)=0$. Consider the vector field $f(x)\frac{\partial}{\partial x}$ on $[0,1]\times[0,1]$, where $x$ is the first coordinate. We can then form a continuous flow $\phi_{t}:B[M,\operatorname{Hur}^{X},N]\times[0,\infty)\to B[M,\operatorname{Hur}^{X},N]$ by moving points in a configuration along the flow, and having them act on the elements $(m,n)$ in $M\times N$ when they reach the boundary. We can define such a flow (which we give the same name) on $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ via the same procedure.
Proposition A.4.0.
Proof.
$Q\to B[M,\operatorname{Hur}^{X},N]$
We define the interior $Q^{\circ}$ to be the dense subset of $Q$ on which the configurations are contained in $(0,1)\times[\epsilon,1-\epsilon]$. In other words, a point of $Q^{\circ}$ is given by $m\in M,n\in N$, such a configuration $x$ and a function $l:x\to X$. We define a continuous function $g:Q^{\circ}\to B[M,\operatorname{Hur}^{X},N]$ as we describe next.
We first continuously choose a homotopy class of paths $\gamma$ from the configuration to $\mathrm{bigord}$ as follows: given a configuration, because the projection $\pi_{2}$ to the second coordinate is injective, we can linearly move all points so that their first coordinate is $\frac{1}{2}$. Then, we can rotate the configuration in $[0,1]\times[0,1]$ counterclockwise by $\frac{\pi}{2}$ so that the second coordinate of all the points is $\frac{1}{2}$, so that the path ends in $\mathrm{bigord}$. This ends our description of $\gamma$.
We claim the above map $Q^{\circ}\to B[M,\operatorname{Hur}^{X},N]$ continuously extends to a map $Q\to B[M,\operatorname{Hur}^{X},N]$. To see this, we can produce a continuous inward flow $\varphi_{-t}$ on $Q$ by pushing points inwards along the vector field $-f(x)\frac{\partial}{\partial x}$ as in § A.4. For positive $t$, this flow lands inside $Q^{\circ}$, so we can continuously extend the function $g$ to a map $\tilde{g}:Q\to B[M,\operatorname{Hur}^{X},N]$ via the formula $\phi_{t}\circ g\circ\varphi_{-t}$ for any $t>0$, with $\phi_{t}$ as in § A.4. It is easy to see that the underlying configuration associated to the map $\tilde{g}$ is similar to the configuration of the original point, with the only difference being that all of the points on the boundary of $[0,1]\times[0,1]$ are removed.
We now establish relation $(ii)$ from § A.4 in this case, as $(i)$ follows from a similar argument. Suppose that we have a point $w\in Q$ where $y$ is the only point in the configuration associated to $w$ on the boundary of $[0,1]\times[0,1]$. Moreover, assume $y$ lies on $1\times[0,1]$. We need to compute the limit of $\phi_{s}\circ g\circ\varphi_{-t}(w)$ as $s$ approaches $t$ from below. For the configuration associated to point $\varphi_{-t}(w)$, we can choose some $\delta<1$ so that the point $y^{\prime}$ to which $y$ flows in in $\varphi_{-t}(w)$ has $\pi_{1}(y^{\prime})>\delta$ (where $\pi_{1}(y^{\prime})$ denotes the first coordinate of $y^{\prime}$) and is the only point of the configuration with first coordinate $\geq\delta$. Let $[\gamma]$ be the homotopy class of paths from the configuration in $\varphi_{-t}(w)$ to $\mathrm{bigord}$ that is used in defining the map $g$. We can choose a different homotopy class of paths by first cutting up $[0,1]\times[0,1]$ into $[0,\delta]\times[0,1]$ and $[\delta,1]\times[0,1]$, doing the construction as in the map $g$ on each piece [^6] Really, the rotation used in the construction should be rescaled since $[0,\delta]\times[0,1]$ and $[\delta,1]\times[0,1]$ are not squares., and concatenating to get a new homotopy class of paths $[\gamma^{\prime}]$. One checks that the class $[\gamma^{\prime}][\gamma]^{-1}$ in the braid group is given by $\sigma_{k-1}\cdots\sigma_{i}$, where there are $k$ points of the configuration, and where $y$ has the $i$th highest second coordinate of all elements in the configuration. We have illustrated checking this in Figure 3. Thus $g\circ\varphi_{-t}(w)$ is equivalent to the point where the path $[\gamma^{\prime}]$ is used instead of $[\gamma]$, and the element of $X^{k}$ is changed according to the action of $\sigma_{k-1}\cdots\sigma_{i}$. Letting $s$ approach $t$, we see that this is exactly the image of the point as described in relation $(ii)$.
@@
Figure 2. This is a picture of the scanning argument of § A.4. In the diagrams, there vertical distance at least $\varepsilon$ between any two red points, and all red points have second coordinate lie in $[\epsilon,1-\epsilon]$.
Figure 3. This is a picture depicting the compatibility of the map $\tilde{g}$ in the proof of § A.4 with relations $(i)$ and $(ii)$ from § A.4. The first two columns depict the paths $\gamma$ and $\gamma^{\prime}$ the points take during the map $\tilde{g}$ under relation $(i)$, and the composite $\gamma^{\prime}\gamma^{-1}$ is depicted underneath. The last two columns depict the paths $\gamma$ and $\gamma^{\prime}$ which the points take under relation $(ii)$, and the composite $\gamma^{\prime}\gamma^{-1}$ is again depicted underneath. In the diagrams, there are $4$ points, $p_{1}$ which is purple, $p_{2}$ which is red, $p_{3}$ which is green, and $p_{4}$ which is blue. So $r=2$ and $k=4$ in the notation of § A.4. The point $p_{2}$ is on the boundary. When $p_{2}$ is on the left boundary, the path is $\sigma_{1}\cdots\sigma_{r-1}=\sigma_{1}$ since $r=2$. When $p_{2}$ is on the right boundary, the path is $\sigma_{r}\cdots\sigma_{k-1}=\sigma_{2}\sigma_{3}$.
@@
The inclusions $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]\to B[M,\operatorname{Hur}^{X},N]$ of § A.4 indexed over the poset of real numbers $01>\epsilon>$ 0 are an ind-weak equivalence.
@@
with $K,K^{\prime}$ compact. Choose a $\delta,\epsilon>0$ such that for all points in the image of $f:K\to B[M,\operatorname{Hur}^{X},N]$, the part of the configuration in $[\frac{1}{2}-\delta,\frac{1}{2}+\delta]\times[0,1]$ projects injectively under the second projection and lands in $[\epsilon,1-\epsilon]$, and each two points in the image of this projection have distance at least $\epsilon$ from each other. Since the flow in § A.4 sends $\frac{1}{2}\pm\delta$ to the boundary in a finite amount of time, we can use the flow to produce a continuous homotopy starting from $f$ and ending with a map landing in the image of $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. First, we verify condition (2) from § A.4. Indeed, since $K$ is compact, at time $t=1$ of the homotopy, the map satisfies the criterion of the last sentence of § A.4, so factors through $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. Finally, we verify condition $(1)$ of § A.4 by using the flow on $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. Indeed, the homotopy also restricts on $K^{\prime}$ to a homotopy of pointed maps to $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. This verifies the conditions of § A.4, as desired. ∎
@@
It is easy to see that for each $01>\epsilon>$ 0, $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ admits the structure of a CW-complex of dimension $<\frac{2}{\epsilon}$. This is because each component of $Q_{\epsilon}[M,\operatorname{Hur}^{X},N]$ is a union of discs, and the gluing relations to form $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ are along subcomplexes. It is finite dimensional since $\epsilon$ puts a bound on the number of points in the configurations that are allowed. For a suitably chosen CW structure on $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$, the inclusion $A\subset\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ is that of a subcomplex, showing the lemma. ∎
@@
$\overline{Q}^{*}_{\epsilon}[M,\operatorname{Hur}_{+}^{X},N]:=\overline{Q}_{\epsilon}[M,\operatorname{Hur}_{+}^{X},N]/S_{M,X,N},$
as $\epsilon$ approaches 00 (for $10<\epsilon<$ 1) where $S_{M,X,N}\subset\overline{Q}_{\epsilon}[M,\operatorname{Hur}_{+}^{X},N]$ is the subspace consisting of all points whose projection to either $M$ or $N$ is the base point.
@@
(A.4) ${*\otimes_{\operatorname{Hur}^{X}}*}$${M\otimes_{\operatorname{Hur}^{X}}*}$${*\otimes_{\operatorname{Hur}^{X}}N}$${M\otimes_{\operatorname{Hur}^{X}}N.}$
@@
are cofibrations, i.e., they both have the homotopy extension property. Hence, the total cofiber of the square

### v3

@@ A.4. A simple topological model
We finish the proof by showing that show that $\operatorname{colim}_{\alpha\in P}\pi_{0}\Omega X^{\prime}_{\alpha}\to\operatorname{colim}_{\alpha\in P}\pi_{0}\Omega X_{\alpha}$ is a bijection. We first check it is surjective. Given a point $\in\Omega X_{\alpha}$, taking $K$ to be the compact space $_{+}$ and $K^{\prime}\subset K$ the base point, hypotheses $(2)$ applied to the pointed map $_{+}\to\Omega X_{\alpha}$ shows surjectivity. For injectivity, if $0,1]_{+}\to\Omega X_{\alpha}$ is a path between two points with the end points factoring through $\Omega X^{\prime}_{\alpha}$, then applying hypothesis $(1)$ to this map, we get a map $0,1]_{+}\wedge[0,1]_{+}=([0,1]\times[0,1])_{+}\to\Omega X_{\beta}$, and the part of the boundary $\{0,1\}\times[0,1]\cup[0,1]\times\{1\}$ yields a path in $\Omega X^{\prime}_{\beta}$ between the two points. ∎
@@
Figure 3. This is a picture of the scanning argument of § A.4. In the diagrams, there vertical distance at least $\varepsilon$ between any two red points, and all red points have second coordinate lie in $\epsilon,1-\epsilon]$.
@@
The inclusions $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]\to B[M,\operatorname{Hur}^{X},N]$ of § A.4 indexed over the poset of real numbers $1>\epsilon>0$ are an ind-weak equivalence.
@@
with $K,K^{\prime}$ compact. Choose a $\delta,\epsilon>0$ such that for all points in the image of $f:K\to B[M,\operatorname{Hur}^{X},N]$, the part of the configuration in $\frac{1}{2}-\delta,\frac{1}{2}+\delta]\times[0,1]$ projects injectively under the second projection and lands in $\epsilon,1-\epsilon]$, and each two points in the image of this projection have distance at least $\epsilon$ from each other. Since the flow in § A.4 sends $\frac{1}{2}\pm\delta$ to the boundary in a finite amount of time, we can use the flow to produce a continuous homotopy starting from $f$ and ending with a map landing in the image of $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$.
We conclude by verifying the two conditions of § A.4. First, we verify condition (2) from § A.4. Indeed, since $K$ is compact, at time $t=1$ of the homotopy, the map satisfies the criterion of the last sentence of § A.4, so factors through $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. Finally, we verify condition $(1)$ of § A.4 by using the flow on $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. Indeed, the homotopy also restricts on $K^{\prime}$ to a homotopy of pointed maps to $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. This verifies the conditions of § A.4, as desired. ∎
@@
It is easy to see that for each $1>\epsilon>0$, $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ admits the structure of a CW-complex of dimension $\frac{2}{\epsilon}$. This is because each component of $Q_{\epsilon}[M,\operatorname{Hur}^{X},N]$ is a union of discs, and the gluing relations to form $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ are along subcomplexes. It is finite dimensional since $\epsilon$ puts a bound on the number of points in the configurations that are allowed. For a suitably chosen CW structure on $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$, the inclusion $A\subset\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ is that of a subcomplex, showing the lemma. ∎
@@
$\overline{Q}^{*}_{\epsilon}[M,\operatorname{Hur}_{+}^{X},N]:=\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]/S_{M,X,N},$
as $\epsilon$ approaches 0 (for $0<\epsilon<1$) where $S_{M,X,N}\subset\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ is the subspace consisting of all points whose projection to either $M$ or $N$ is the base point.
@@
(A.4) $\otimes_{\operatorname{Hur}^{X}}*$$M\otimes_{\operatorname{Hur}^{X}}*$$\otimes_{\operatorname{Hur}^{X}}N$$M\otimes_{\operatorname{Hur}^{X}}N.$}
@@
are cofibrations, i.e., they both have the homotopy extension property. In turn, this implies the map
@@
is a cofibration, where the $\cup$ notation denotes the pushout. Finally, this implies the total cofiber of the square